OpenStudy (anonymous):

Multiply and simplify: x + 11 over x + 3 times x squared minus 9 over x squared + 20x + 99

OpenStudy (aroub):

x+11 x^2-9 = ---- * ------ x+3 x^2+20x+99 factor, x+11 (x-3)(x+3) = ---- * ----------- x+3 (x+9)(x+11) cancel, (x-3) = ---- (x+9)

OpenStudy (anonymous):

thank you for explaining

OpenStudy (aroub):

np ! and thanks for the medal :)

OpenStudy (anonymous):

no problem